Grade 12Algebra

Replace the blank with an integer to make it a true statement.
5 × _____ = –35

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us assume the missing integer be x,
Then,
= (5) × (x) = -35
= x = - (-35/5)
= x = -7
Let us substitute the value of x in the place of blank,
= (5) × (-7) = -35 ... [∵(+ × - = -)]

Thanks